# A Hybrid Transformer-Mamba Language Model

### v1

@@ 1 Introduction
We introduce Jamba, a new publicly available large language model. Jamba is based on a novel hybrid architecture, which combines Transformer layers [46] with Mamba layers [16], a recent state-space model [18, 17], as well as a mixture-of-experts (MoE) component [41, 13]. Jamba thus combines two orthogonal architectural designs that together give it improved performance and higher throughput, while maintaining a manageable memory footprint. The 7B-based Jamba model (12B active parameters, 52B total available parameters) we are releasing was designed to fit in a single 80GB GPU, but the Jamba architecture supports other design choices, depending on one’s hardware and performance requirements.
@@
Somewhat unusually for a new architecture, we release Jamba (12B active parameters, 52B total available parameters) under Apache 2.0 license: https://huggingface.co/ai21labs/Jamba-v0.1. We do so since we feel that the novel architecture of Jamba calls for further study, experimentation, and optimization by the community. Our design was based on various ablation experiments we conducted to explore the effect of different tradeoffs and design choices, and insights gleaned from those. These ablations were performed at scales of up to 7B parameters, and training runs of up to 250B tokens. We plan to release model checkpoints from these runs.
@@ 5.1 Academic Benchmarks
Noticebly, Jamba performs comparably to the leading publicly available models of similar or larger size, including Llama-2 70B and Mixtral. At the same time, our model has a smaller number of total available parameters than Llama-2 (52B compared to 70B). Moreover, as a sparse model, Jamba has only 12B active parameters, similar to Mixtral’s 12.9B active parameters. However, as a fully-attentional model, Mixtral has a large memory footprint with long sequences, requiring 32GB for the KV cache with 256K tokens. In contrast, thanks to its hybrid Attention-Mamba architecture, Jamba’s KV cache takes only 4GB even at such a long context (Section 2). Importantly, our Jamba achieves such a strong performance while having much better throughput than Llama-2 70B and Mixtral, up to 3x improvement (Section 3.2).
In summary, Jamba demostrates the ability of hybrid architectures to reach the performance of state-of-the-art Transformer based models of the same size class, while having the benefits of an SSM.
@@
Table 2: Comparison of Jamba with other publicly available models. Jamba obtains similar or better performance with much better throughput.
@@ 5.2 Long-Context Evaluations
We have successfully trained Jamba models with context lengths of up to 1M tokens. The released model handles context lengths of up to 256K tokens. In this section, we evaluate it on synthetic and naturalistic benchmarks that test is long-context capabilities.
@@ 5.2.2 Naturalistic long-context evaluation
We evaluate Jamba’s ability to handle long contexts using question-answering benchmarks, consisting of long inputs. To this end, we repurpose five of the longest-context datasets from L-Eval [2], by structuring them in a few-shot format (we use 3-shots in all experiments here). Specifically, we evaluated the models on the following datasets: NarrativeQA (QA on narratives; [25]), LongFQA (finance; [2]), Natural Questions (NQ; Wikipedia; [26]), CUAD (law; [21]), and SFiction (science fiction). The average input length in these datasets ranges from 6K to 62K tokens. These context lengths are further highly expanded by the few-shot format.
Table 3 summarizes the evaluation results, in terms of F1. Jamba outperforms Mixtral on most of the datasets as well as on average. In addition, as these long-context tasks require substantial computation, here Jamba’s efficiency shines, with much better throughput with long contexts (Section 3.2).
@@ 6 Ablations and Insights
This section discusses ablation experiments we ran for different design choices in our implementation of the Jamba architecture. First we show the benefit of combining attention and Mamba layers, at which ratio they should be combined, and how to interleave them. We investigate cases where pure Mamba fails, suggesting that it struggles to develop in-context learning capabilities, while the Attention–Mamba hybrid exhibits in-context learning similar to vanilla Transformers. Then we show the benefit of adding MoE on top of a hybrid Attention–Mamba model. Finally, we share two additional learnings that we found useful: explicit positional information is not needed in Jamba, and Mamba layers necessitate special normalization to stabilize training at large scale. [^3] In all the ablation experiments, “pure Mamba” refers to models with Mamba layers interleaved with MLP layers.
@@ 6.1 Benefits of combining Attention and Mamba
Figure 5: Training loss curves for pure Attention, pure Mamba, and Attention-Mamba hybrids (no MoE), with ratios $a:m$ of 1:3 and 1:4. All models are 1.3B parameters. The two hybrids achieve better loss throughout this training run, without any noticeable difference between the different Attention/Mamba ratios.
Next, we compare performance of vanilla Transformer, vanilla Mamba, and Attention-Mamba hybrid models, at 7B model size, after training on 50B tokens. As Table 5 shows, the pure Mamba layer is quite competitive, but lags slightly behind pure Attention. The hybrid Attention-Mamba (without MoE) outperforms the pure models while obtaining better throughput than the vanilla Transformer (Section 3.2).
@@
Figure 6: Training loss curves for pure Attention, pure Mamba, and an Attention-Mamba hybrid (no MoE). All models are 7B parameters. the hybrid achives better loss throughout this training run.

### v2

@@ 1 Introduction
We introduce Jamba, a new publicly available large language model. Jamba is based on a novel hybrid architecture, which combines Transformer layers [51] with Mamba layers [17], a recent state-space model [19, 18], as well as a mixture-of-experts (MoE) module [46, 14]. Jamba thus combines two orthogonal architectural designs that together give it improved performance and higher throughput, while maintaining a manageable memory footprint. The 7B-based Jamba model (12B active parameters, 52B total available parameters) we are releasing was designed to fit in a single 80GB GPU, but the Jamba architecture supports other design choices, depending on one’s hardware and performance requirements.
@@
Somewhat unusual for a new architecture, we release Jamba (12B active parameters, 52B total available parameters) under Apache 2.0 license: https://huggingface.co/ai21labs/Jamba-v0.1. We do so since we feel that the novel architecture of Jamba calls for further study, experimentation, and optimization by the community. Our design was based on various ablation experiments we conducted to explore the effect of different tradeoffs and design choices, and insights gleaned from those. These ablations were performed at scales of up to 7B parameters, and training runs of up to 250B tokens. We plan to release model checkpoints from these runs.
@@ 5.1 Academic Benchmarks
In most tasks, Jamba performs comparably to leading publicly available models of similar or larger size, including Llama-2 70B and Mixtral. At the same time, our model has a smaller number of total available parameters than Llama-2 (52B compared to 70B). Moreover, as a sparse model, Jamba has only 12B active parameters, similar to Mixtral’s 12.9B active parameters. However, as a fully-attentional model, Mixtral has a large memory footprint with long sequences, requiring 32GB for the KV cache with 256K tokens. In contrast, thanks to its hybrid Attention-Mamba architecture, Jamba’s KV cache takes only 4GB even at such a long context (Section 2). Importantly, our Jamba achieves such a strong performance while having much better throughput than Llama-2 70B and Mixtral, up to 3x improvement (Section 3.2).
In summary, Jamba demostrates the ability of hybrid architectures to reach the performance of state-of-the-art Transformer models of the same size class, while having the benefits of an SSM.
@@
Table 2: Comparison of Jamba with other publicly available models. Jamba obtains similar performance with much better throughput.
@@ 5.2 Long-Context Evaluations
We have successfully trained Jamba models with context lengths of up to 1M tokens. The released model handles context lengths of up to 256K tokens. In this section, we evaluate it on synthetic and naturalistic benchmarks that test its long-context capabilities.
@@ 5.2.2 Naturalistic long-context evaluation
We evaluate Jamba’s ability to handle long contexts in two settings. First, we evaluate the model on several classification tasks commonly used for assessing in-context learning, with an increasing number of few-shot examples. In particular, we use the four datasets with the largest label space from [41], which showed that such tasks benefit most from using more few-shot examples: Trec-Fine (fine-grained question type classification, 50 labels; [29]), NLU Intent (intent classification in natural language understanding, 68 labels; [31]), Banking77 (intent classification in the banking domain, 77 labels; [4]), and CLINC150 (intent classification, 150 labels; [28]). In each case we add examples up to a context length of 128K tokens. Figure 5 shows the results, comparing Jamba to Mixtral. In Trec-Fine and Banking77, Jamba outperforms Mixtral, especially with a large number of few-shot examples. In NLU Intent and CLINC150, the two models are on par.
(a) TREC Fine-grained.
(b) NLU Intent.
(c) Banking77.
(d) CLINC150.
Figure 5: Comparison of Jamba and Mixtral on few-shot classification with a large number of few-shot examples. Results are exact match with greedy decoding.
Second, we use question-answering datasets, consisting of long inputs. To this end, we repurpose five of the longest-context datasets from L-Eval [2], by structuring them in a few-shot format (we use 3-shots in these experiments). Specifically, we evaluated the models on the following datasets: NarrativeQA (QA on narratives; [26]), LongFQA (finance; [2]), Natural Questions (NQ; Wikipedia; [27]), CUAD (law; [22]), and SFiction (science fiction). The average input length in these datasets ranges from 6K to 62K tokens. These lengths are further highly expanded by the few-shot format.
Table 3 summarizes the results, in terms of F1.333F1 score is the recommended metric in L-Eval [2]. In addition, our setup calibrates the length of the few-shot completions to approximately match the length of the test completion, increasing the credibility of F1. Jamba outperforms Mixtral on most of the datasets as well as on average. In addition, as these long-context tasks require substantial computation, here Jamba’s efficiency shines, with much better throughput with long contexts (Section 3.2).
@@ 6 Ablations and Insights
This section discusses ablation experiments we ran for different design choices in our implementation of the Jamba architecture. First we show the benefit of combining attention and Mamba layers, at which ratio they should be combined, and how to interleave them. We investigate cases where pure Mamba fails, suggesting that it struggles to develop in-context learning capabilities, while the Attention-Mamba hybrid exhibits in-context learning similar to vanilla Transformers. Then we show the benefit of adding MoE on top of a hybrid Attention-Mamba model. Finally, we share two additional learnings that we found useful: explicit positional information is not needed in Jamba, and Mamba layers necessitate special normalization to stabilize training at large scale. [^4] In all the ablations, “pure Mamba” refers to models with Mamba layers interleaved with MLP layers.
@@ 6.1 Benefits of combining Attention and Mamba
Figure 6: Training loss curves for pure Attention, pure Mamba, and Attention-Mamba hybrids (no MoE), with ratios $a:m$ of 1:3 and 1:7. All models are 1.3B parameters. The two hybrids achieve better loss throughout this training run, without any noticeable difference between the different Attention/Mamba ratios.
Next, we compare performance of vanilla Transformer, vanilla Mamba, and Attention-Mamba hybrid models, at 7B model size, after training on 50B tokens. As Table 5 shows, the pure Mamba model is quite competitive, but lags slightly behind pure Attention. The hybrid Attention-Mamba (without MoE) outperforms the pure models while obtaining better throughput than the vanilla Transformer (Section 3.2).
@@
Figure 7: Training loss curves for pure Attention, pure Mamba, and an Attention-Mamba hybrid (no MoE). All models are 7B parameters. The hybrid achieves better loss throughout this training run.